Draw a pedigree for at least three generations of a family where an X-linked dominant trait is being inherited. Include an affected male and an affected female in your pedigree as well as unaffected individuals. Write out the genotypes for an affected male, affected female, unaffected male, and unaffected female.


1
Expert's answer
2021-09-29T00:48:02-0400


Affected male "= X^AY"

affected female "= X^aX^A , X^AX^a"

unaffected male "= X^aY"

unaffected female "= X^aX^a"